### v1

# Framing Unionization on Facebook: Communication around Representation Elections in the United States

Abstract
Digital media have become central to how labor unions communicate, organize, and sustain collective action. Yet little is known about how unions’ online discourse relates to concrete outcomes such as representation elections. This study addresses the gap by combining National Labor Relations Board (NLRB) election data with 158k Facebook posts published by U.S. labor unions between 2015 and 2024. We focused on five discourse frames widely recognized in labor and social movement communication research: diagnostic (identifying problems), prognostic (proposing solutions), motivational (mobilizing action), community (emphasizing solidarity), and engagement (promoting interaction). Using a fine-tuned RoBERTa classifier, we systematically annotated unions’ posts and analyzed patterns of frame usage around election events. Our findings showed that diagnostic and community frames dominated union communication overall, but that frame usage varied substantially across organizations. In election cases that unions won, communication leading up to the vote showed an increased use of diagnostic, prognostic, and community frames, followed by a reduction in prognostic and motivational framing after the event—patterns consistent with strategic preparation. By contrast, in lost election cases unions showed little adjustment in their communication, suggesting an absence of tailored communication strategies. By examining variation in message-level framing, the study highlights how communication strategies adapt to organizational contexts, contributing open tools and data and complementing prior research in understanding digital communication of unions and social movements.
Introduction
Digital media has transformed the landscape of collective action. Civil Society Organizations (CSOs) and social movements increasingly rely on online platforms to mobilize participants, coordinate activities, and frame issues to reach broader audiences. Scholars have highlighted the promises of these digital technologies: lowering barriers to participation, enabling new forms of connective action, and amplifying marginalized voices. At the same time, they have cautioned about their limitations, including resource constraints, fragmented attention, and uncertain effects on long-term organizational capacity (Bennett and Segerberg 2012; Earl and Kimport 2011; Tufekci 2017).
Labor unions provide a distinctive case within the ecosystem of CSOs and social movements. As one of the most institutionalized forms of collective organization, they have experienced declining membership and political influence in the United States since the mid-twentieth century. This led scholars to ask whether their revitalization requires not only traditional collective bargaining but also new forms of communication and public engagement (Milkman 2013; Hurd 1998; Voss and Sherman 2000). Social media platforms offer this type of strategic opportunity, especially in today’s fragmented workforce with precarious employment and immigrant labor (Milkman 2020). Unions have indeed turned to digital media as a strategic opportunity: a tool to communicate with members, reach new audiences, and sustain organizational relevance (Lee 1996). Among digital platforms, Facebook has emerged as particularly important (Carneiro and Costa 2022; Ford and Sinpeng 2022). Like other social media, it supports both outward and inward communication, and attracts users whose demographic profiles resemble many unionized workforces. These features make it strategically relevant for labor mobilization (Blanc 2022). Despite this promise, the platform also introduces new complexities, such as uncertainty about which engagement strategies are most effective (Greene and Kirton 2003), and little is known about how union communication on the platform relates to concrete unionization outcomes at scale, particularly in representation elections.
In this study, we combined NLRB election data with a large corpus of Facebook posts by labor unions in the United States from 2015 to 2024 to examine the relationship between online communication and election outcomes. Using natural language processing and computational methods, we analyzed how unions framed their digital communication around representation elections and whether these patterns varied by electoral outcomes [^1] Data and code are available at
https://github.com/ariannap13/laborunions. Models are available on HuggingFace as ariannap22/roberta˙finetuned˙fb˙unions and ariannap22/roberta˙finetuned˙multilabel˙unions
. Specifically, we addressed three research questions:
RQ1. To what extent do labor unions use discourse frames in their Facebook communication?
RQ2. Does the use of such frames leading up to an election differ for won and lost cases?
RQ3. How do patterns of frame usage change after election events, and do these dynamics vary between wins and losses?
Our contributions are threefold: (i) we provide guidelines for annotating discourse frames in unions’ social media text, with possible application to organizations beyond the union context and to platforms beyond Facebook; (ii) we release a fine-tuned, ready-to-use classification model for detecting frames in unions’ communication; and (iii) we advance research at the intersection of framing and labor movement, offering new insights into the relationship between digital communication and unionization dynamics. While much labor movement scholarship often treats unions at the collective level  (Snow, Benford et al. 1988; Benford and Snow 2000; Kelly 1998; Voss and Sherman 2000), we show how digital platforms enable the observation of framing strategies at the level of individual messages and connect them to concrete organizational outcomes. This micro-comparative approach suits well the heterogeneity of communication practices across unions, and allows us to quantify the role of framing as a mechanism of adaptation in periods of organizational change.
Building on this perspective, our analysis of union communication on Facebook revealed several key patterns, many of which are in line with prior literature. We found that unions predominantly emphasized problem identification and community-building, though frame use varied across organizations. The mobilization potential, embodied by a discourse centered on call-to-action messages, remained largely untapped. These findings align with research on union and civil society organizations on social media (Lovejoy and Saxton 2012; Uba and Jansson 2021; Carneiro and Costa 2022). Unions representing broad, multi-industry workforces (industrial unions) tended to emphasize frames for grievances and solutions, whereas unions representing specialized trades (craft unions) relied less on these frames. Prior to elections, in cases that unions ultimately won, communication intensified around problem identification, solution-oriented messaging, and community building, while solution-oriented and call-to-action frames generally declined after the election. In cases that unions lost, by contrast, framing patterns remained relatively stable from pre- to post-election periods, suggesting that strategic adjustments in communication may play a role in shaping election outcomes. This link between communication strategies and organizational outcomes, as well as their adjustment after victory, is also reflected in prior work on union revitalization (Turner 2005) and theories of organizational life cycles of social movements (Christiansen 2009; Tarrow 2022).
Our findings provide actionable insights for labor unions, civil society organizations, and social movements on their digital communication strategies. In particular, using informal platforms such as social media to discuss issues and propose solutions may be more effective than relying on these channels for superficial engagement, which often dominates professionalized social media management. Beyond these practical implications, our results contribute to theoretical debates on framing and collective action by showing how message-level strategies vary across organizations and electoral contexts. By shifting focus from collective-level analyses of movements to the micro-dynamics of discourse on digital platforms, this study highlights how framing might operate as a mechanism of adaptation in periods of organizational change.
Theoretical Framework: Discourse Frames
Labor unions, like other civil society organizations and social movements, rely on communication to achieve core organizational goals. Nowadays much of this activity occurs in digital spaces, making it critical to understand the communicative strategies unions employ online.
The theory of collective action frames (Snow, Benford et al. 1988; Benford and Snow 2000) offers a central lens for analyzing these strategies. It identifies the following “schemata of interpretation” (Goffman 1974), or functions of social movements’ communication: diagnostic frames, which highlight problems and assign responsibility; prognostic frames, which propose solutions; and motivational frames, which seek to mobilize audiences. These three core frames also parallel the conditions for union revitalization outlined by Kelly (1998): workers must recognize grievances (diagnostic), perceive them as actionable (prognostic), and see unions as legitimate vehicles for change (motivational).
To capture the specificity of union communication online, we extended this framework with two additional categories. First, community frames that emphasize solidarity, shared identity, and collective strength—aspects particularly salient in union contexts (Carneiro and Costa 2022) and consistent with the community cultivation function in nonprofit communication (Lovejoy and Saxton 2012). Second, engagement frames that capture efforts to stimulate interaction, circulation, and participation, reflecting the affordances and dynamics of social media (Muntinga, Moorman, and Smit 2011).
In summary, our analysis operationalized five discourse frames:
  • Diagnostic: identifying problems and attributing responsibility.
  • Prognostic: advancing solutions and strategies for change.
  • Motivational: encouraging participation through calls to action.
  • Community: fostering solidarity and collective identity.
  • Engagement: promoting interaction and audience involvement.
This framework maps prior research onto our operationalization, providing a theoretically grounded basis for analyzing union social media communication.
Methodological Framework
Data Curation
Our focus is on the central stage of the unionization process: representation elections and their outcomes (win or loss). We collected data on election results from publicly available records of the National Labour Relations Board (NLRB) (National Labor Relations Board 2025), covering all cases for which a representation election (RC cases in the NLRB data) was petitioned and held between January 2015 and December 2024 for labor unions in the United States. From the available variables, we focused on the numeric identifier of each election case, the election date, the names of participating unions, and the union that won the election.
The dataset comprises 11,416 election cases, of which 10,975 involve a single union, 417 involve two unions, and 24 involve three unions competing in the same election. Since much of the information was entered manually in the NLRB data, union names exhibit considerable variation in formatting and granularity. To facilitate analysis of unions’ Facebook (FB) activity, we standardized these names at the national or international level. For example, “Communications Workers of America Local 7250” was mapped to “Communications Workers of America.” To identify a set of national and international unions, we began with the official list of unions affiliated with the American Federation of Labor and Congress of Industrial Organizations (AFL-CIO) [^2] AFL-CIO affiliated unions. https://aflcio.org/about-us/our-unions-and-allies/our-affiliated-unions. Accessed March 3, 2025., consolidating subunits under their parent organizations (e.g., TNG-CWA under CWA) using regular expressions and rule-based assignments. This process yielded 47 unions. For unions not included in the AFL-CIO list, we identified common patterns and acronyms using a combination of regular expressions (e.g., extracting acronyms in parentheses) and manual grouping, resulting in additional 510 potential national and international unions.
To link these organizations with their digital presence, we queried the Bing Search API [^3] Microsoft Bing Search API. https://api.bing.microsoft.com/v7.0/search. Accessed March 3, 2025 with each union name alongside the string “union Facebook” to identify official FB pages. We parsed the results using regular expressions and manually verified their accuracy. This resulted in 66 unions with official FB accounts, of which 40 provided data through the Meta Research API Content Library (Meta Platforms, Inc. 2024). Table D1 lists the unions involved in the study, together with their acronyms and type of structure. In total, we collected 158,238 posts (157,883 of which had text available) published between January 2015 and December 2024, matching the time span of the election data.
Of the 11,416 NLRB election cases, we identified 2,607 instances where unions with available FB data lost and 6,445 where they won.
Discourse Frames Classifier
We trained a classifier to detect five discourse frames in union Facebook posts: diagnostic, prognostic, motivational, community, and engagement. Aiming at capturing the nuanced usage of such frames in online communication, we set the task to be multi-label. Following common practices in Natural Language Processing (NLP), we first fine-tuned a RoBERTa-base model (Liu et al. 2019) on the text of the 157,883 posts to better capture typical linguistic patterns of labor union communication on Facebook. After collecting annotations for the presence of discourse frames [^4] We also annotated posts for an additional political endorsement label to capture explicit support for political actors, but given its low prevalence, it was not included in the main analysis; see Appendix A and B., we further fine-tuned this model to perform the multi-label classification task. Details on model configuration and computational resources are provided in Appendix B.
Frames Annotation
We conducted a manual annotation campaign with two of the authors as coders, considering a multi-label setting for the presence of discourse frames. Starting with an initial set of 100 posts and taking inspiration from prior work on measuring collective action frames in text (Fernández-Zubieta et al. 2023; Mendelsohn et al. 2024), we iteratively defined a codebook for the identification of discourse frames of interest. The codebook is reported in Appendix A.
Using this codebook, the coders annotated a new sample of 200 posts, achieving a Cohen’s $\kappa$ inter-annotator agreement of 0.8 across labels. Given this high level of agreement, the coders then independently annotated disjoint sets of additional randomly sampled posts. In total, this resulted in 1,151 new posts, each annotated by a single coder.
Multi-label Classification
We randomly split the annotated dataset into training (80%) and test (20%) sets. The RoBERTa model pre-adapted to Facebook language was then fine-tuned on the training set for the multi-label classification task. After evaluating its performance on the test set, we applied the final classifier to the entire corpus to generate frame predictions for all posts.
Patterns of Frames Usage
We examined how unions’ usage of discourse frames differed in election cases that were won versus lost in the days preceding the vote, and how these patterns changed afterward.
To account for sparsity in posting frequency, we aggregated our measurements using a rolling window of $n_{days}=5$. Given an active union on Facebook, for each day and each frame, we calculated (i) the number of posts containing that frame and (ii) the proportion of such posts relative to all posts on that day. To focus on activity around elections, for each election case, we defined two non-overlapping windows:
  • Pre-election: mean proportion of posts expressing a given frame from 7 to 3 days before the election.
  • Post-election: mean proportion from 3 to 7 days after the election.
These windows exclude the election day itself—avoiding overlap caused by the rolling aggregation—while capturing both the immediate lead-up and aftermath.
Detrended Frames Usage and its Change
Observed frame usage patterns may be driven by long-term temporal dynamics rather than immediate responses to specific events (e.g., a gradual increase in certain frames over time). To separate event-specific effects from such background trends, we computed detrended scores. For each union, we identified baseline periods of 18 consecutive days during the reference period (January 2015–December 2024) in which the union was not involved in a specific election outcome (win or loss). We assigned a pseudo-event date at the midpoint of each baseline period and applied the same pre- and post-event aggregation procedure as for real elections.
We then computed the detrended usage of each frame as
${U_{b}}^{D}=U_{b}-{U_{b}}^{baseline},{U_{a}}^{D}=U_{a}-{U_{a}}^{baseline}$
where $U_{b}$ and $U_{a}$ denote frame usage before and after the election, and ${U_{b}}^{baseline}$ and ${U_{a}}^{baseline}$ are the corresponding baseline values. Values of ${U_{b}}^{D}$ or ${U_{a}}^{D}$ close to 0 indicate frame usage similar to baseline periods, positive values indicate higher-than-baseline usage, and negative values indicate lower-than-baseline usage. By examining pre-election usage separately for cases that were won and cases that were lost, we described how unions adjusted their communication relative to the baseline. Comparing the two groups directly allowed us to identify differences in their pre-election communication strategies.
We classified changes in frame usage from the pre- to post-election periods into three categories: decrease, stable, and increase. For each union–case instance, we computed the change in detrended frame usage as the offset
$O={U_{a}}^{D}-{U_{b}}^{D}$
To assign union-case instances to categories of change, we examined the combined distribution of $O$ values (across both won and lost cases). We then used the 25th and 75th percentiles of this distribution as thresholds:
$O=\begin{cases}\textbf{decrease},&O<\text{percentile}_{25}\\ \textbf{stable},&\text{percentile}_{25}\leq O\leq\text{percentile}_{75}\\ \textbf{increase},&O>\text{percentile}_{75}\end{cases}$
This procedure provided a consistent, data-driven way to categorize changes in frame usage, enabling systematic comparisons between won and lost election cases and across frames.
Robustness Checks
To ensure that results generalize well beyond specific design choices, we conducted several robustness checks: (i) we varied the rolling aggregation window size ($n_{days}=5$ vs $n_{days}=3$) (see Appendix C.1); (ii) we balanced the number of won and lost cases within each labor union, repeating analyses across multiple random seeds (see Appendix C.2); and (iii) we down-sampled highly active unions to the 90th percentile of the distribution of representation cases by labor unions, repeating analyses across multiple random seeds (see Appendix C.2).
Results
Classification of Discourse Frames
We applied our fine-tuned RoBERTa model to Facebook posts to perform multi-label classification of discourse frames. Table 1 reports performance in terms of precision, recall, and F1 scores for the test set. Scores are comparable to similar studies (Fernández-Zubieta et al. 2023; Mendelsohn et al. 2024), with F1 values ranging from 0.68 to 0.78 across frames and a macro F1 of 0.73.
Frame Precision Recall F1-score Support
Diagnostic 0.72 0.85 0.78 84
Prognostic 0.67 0.69 0.68 58
Motivational 0.57 0.83 0.68 29
Community 0.75 0.81 0.78 81
Engagement 0.71 0.78 0.75 51
Micro avg 0.70 0.80 0.74 303
Macro avg 0.68 0.79 0.73 303
Weighted avg 0.70 0.80 0.74 303
Table 1: Performance of the multi-label discourse frames classifier.
We then applied the classifier to the corpus of 157,883 textual posts, identifying 140,199 as expressing at least one of the frames of interest.
Usage of Discourse Frames
Aiming at defining a baseline distribution of frame usage across unions, we first looked at the distributions of posts by union. We identified skewness (minimum = 779, 25th percentile = 1,921, 75th percentile = 5,442, maximum = 10,452, full distribution in Figure D1) and thus implemented a bootstrap-like down-sampling procedure. Over-represented unions were down-sampled to the minimum number of posts in the distribution, and the procedure was repeated with five different random seeds. For each frame, we computed the median proportion of posts expressing that frame across seeds. Overall, the median baseline distribution of frame usage is as follows: diagnostic in 40.9% of posts, community in 39.4%, prognostic in 28.4%, engagement in 23.1%, and motivational in 15.5%.
Figure 1 shows how individual unions diverge from the overall median baseline in their use of different frames. For each union, we calculated the share of posts expressing a given frame relative to its total output, and then expressed this value as a percentage difference over the median baseline. A value of +100% indicates that the union used that frame twice as often as the median baseline, 0% indicates equal usage, and –50% indicates that the union used the frame half as often. Building on a common distinction in labor movement scholarship (Herrigel 1993; Dubofsky 2000), we classified unions into two categories: craft unions, which organize workers sharing a specific trade or skill, and industrial unions, which organize workers across diverse trades, skill levels, and occupations. This classification was determined through manual review of information on unions’ official websites.
Figure 1: Relative usage of discourse frames by labor unions compared to the baseline median (percentage difference over baseline). Red color indicates a usage higher than the baseline, blue color indicates a usage lower than the baseline. The dendrogram shows the clustering of unions based on their framing patterns across the five frame types. Union acronyms are colored according to their organizational structure.
Many unions behaved in line with the baseline; nonetheless, a clear divide emerged. Unions clustered into two broad groups based on whether their use of diagnostic and prognostic frames was higher or lower than the baseline. In particular, industrial unions tended to employ diagnostic and prognostic frames more frequently than the baseline. Examples include the Service Employees International Union (SEIU), the International Federation of Professional and Technical Engineers (IFPTE), the International Brotherhood of Teamsters (IBT), the Amalgamated Transit Union (ATU), National Nurses United (NNU), and Unite Here (UNITE HERE), all located toward the bottom of the dendrogram (union acronyms are expanded in Table D1). By contrast, lower-than-baseline usage of diagnostic and prognostic frames was more characteristic of craft unions. This distinction might reflect different organizational orientations: industrial unions, with their broad membership bases, are more often engaged in large-scale mobilization and bargaining arenas, where diagnostic frames (grievance articulation) and prognostic frames (solution proposals) play a central role (Voss and Sherman 2000). Variation also emerged in the use of the engagement frame. Industrial unions, in particular, often used engagement frames less than the baseline (cf. bottom of the plot), whereas higher-than-baseline engagement showed no consistent relationship to union structure (cf. top of the plot). Notably, the United Union of Roofers, Waterproofers and Allied Workers (Roofers) employed the engagement frame nearly two and a half times more often than the baseline.
Taken together, these results suggest that industrial unions tend to differentiate themselves from craft unions by emphasizing grievance and solution-oriented frames while downplaying interactive engagement, consistent with their larger-scale bargaining and mobilization needs. Craft unions, on the other hand, appear less reliant on diagnostic and prognostic repertoires, and in some cases lean more heavily on engagement, indicating a different balance in communication strategies.
Frames Usage Leading to Elections
We next examined the usage of frames in the lead-up to election events. Figure 2 shows the average proportion of posts that express each frame in the pre-election period and its 95% confidence interval, separately for losing and winning outcomes. The usage of frames has been detrended by considering baseline periods with no election outcomes of a given kind: positive values indicate that a frame was used more often than usual, and negative values indicate lower-than-typical usage. This allowed us to examine how unions changed their frame usage leading up to elections relative to normal periods, and to compare won and lost cases. We down-sampled the number of wins and losses per union to ensure a balanced sample. Results are robust to the choice of sampling seed, as shown in Table C1.
Figure 2: Usage of frames leading to the election, comparing lost and won cases, in terms of mean and 95% confidence interval. Values on the y-axis refer to detrended scores. The $\star$ sign refers to the significance of the test for difference in means between losing and winning: $\star$ indicates significance at $\alpha=0.05$, $\star\star$ at $\alpha=0.01$, and $\star\star\star$ at $\alpha=0.001$.
In both won and lost cases, frame usage shifted before elections, but in different ways. Prior to wins, communication showed increases in all frames relative to baseline, except for engagement, which decreased. Prior to losses, by contrast, usage of motivational and community frames remained close to baseline, while diagnostic, prognostic, and engagement frames slightly decreased.
Next, we compared won and lost representation cases directly. Independent t-tests for the difference between group means showed statistically significant differences for all frames ($\alpha=0.001$, indicated as $\star$$\star$$\star$} in Figure 2), with communication in won cases exhibiting higher usage of nearly all frames compared to lost cases, particularly diagnostic, prognostic, and community frames. Robustness checks using multiple random seeds and down-sampling of over-represented unions showed that the differences in means between losses and wins were consistent for these three frames, while differences for motivational and engagement frames were not always significant (see Appendix Table C2).
Frames Usage Pre- and Post-Elections
We finally examined how frame usage changes from the pre- to the post-election period. We first verified that frame usage before and after elections was robust to the choice of window size for rolling aggregation (see Appendix C.1). For each frame, we classified changes from pre- to post-election into three patterns: decrease, stable, or increase, depending on whether post-election usage was lower, similar, or higher than pre-election usage. We then counted the number of union–case instances with won or lost outcomes assigned to each pattern, providing an overall measure of how prevalent each change pattern was for each frame. We down-sampled the number of wins and losses per union to ensure a balanced sample. Results are robust to the choice of sampling seed, as shown in Figure C3. Figure 3 shows the difference in pattern prevalence between losses and wins, computed as losing minus winning. Positive values indicate a pattern was more common among losses, negative values indicate it was more common among wins. For example, a value of -5% in the decrease pattern for the prognostic frame means that in won cases, frame usage declined 5% more often than in lost cases. We computed Newcombe’s confidence intervals at $\alpha=0.05$ for the difference between the two independent proportions and considered a difference significant if the interval excluded 0.
Figure 3: Changes in frame usage before and after elections, comparing lost and won cases. Bold values in the cells refer to the difference between proportion of pattern presence for a given frame between losses and wins. Raw counts of union-case instances assigned to each cell are reported for losses (L) and wins (W). The $\star$ sign indicates a significant difference between losing and winning at $\alpha=0.05$.
In won cases, frame usage more often decreased for prognostic and motivational frames after elections. In lost cases, by contrast, prognostic and engagement usage more often increased, while motivational usage remained stable.
Robustness checks across multiple random seeds and down-sampling of over-represented unions confirmed that the decreases in prognostic and motivational frames in won cases were consistent. Other patterns were not consistently significant (see Appendix C.2).
Discussion and Conclusion
Traditional sociological research at the intersection of framing and labor movement has often taken a macroscopic perspective when investigating how movement communication relates to organizational outcomes. Such approaches typically conceptualize the labor movement at a collective level. Yet, in the age of connective action (Bennett and Segerberg 2012), political action is increasingly dispersed and decentralized (Kavada 2016), calling for more microscopic perspectives. Within social media contexts, this relates to analyzing how individual messages employ discourse frames and connecting these patterns to concrete organizational outcomes, such as labor unions’ representation elections.
Some qualitative and mixed-methods studies have already demonstrated the value of this kind of granular analysis (Pasquier and Wood 2018; Frangi, Zhang, and Hebdon 2020; Lazar, Ribak, and Davidson 2020). However, these works tend to be limited to single campaigns, making generalization difficult. Quantitative analyses of union communication, in turn, often examine only a restricted number of accounts, or remain disconnected from organizational outcomes (Hodder and Houghton 2020; Houghton and Hodder 2021; Uba and Jansson 2021; Carneiro and Costa 2022). Our study addressed these gaps by contributing a micro-level, comparative analysis of 158k Facebook posts by 40 U.S. labor unions between 2015 and 2024. Building on framing theory applied in labor and social movement research (Kelly 1998; Benford and Snow 2000; Lovejoy and Saxton 2012; Carneiro and Costa 2022), we developed an annotation codebook and a machine learning model to infer the usage of discourse frames, linking them to representation election outcomes. This approach enabled the kind of systematic, comparative work often missing in framing and social movements scholarship (Tarrow 1996; Benford 1997; Snow et al. 2014).
Prior work has documented the untapped mobilization potential of social media for unions and non-profit organizations. Their communication focus is mainly on routine activities, information campaigns, and community-building messages, with very little attention to action-focused messages (Lovejoy and Saxton 2012; Uba and Jansson 2021; Carneiro and Costa 2022). These studies also highlight the diversity of communication strategies employed across organizations. Our findings confirmed and extended these insights.
First, we found that nearly 89% of posts contain at least one frame, with diagnostic and community frames dominating overall discourse (RQ1). This suggests that unions rely heavily on problem identification and community-building to structure their communication. Such practices resonate with the expression of stances on problematic issues, which relate to the diagnostic frame, and celebrations of achievements, common in the community frame, stressed in Carneiro and Costa (2022). Moreover, they are close to the relationship-fostering role of social media messaging identified by Lovejoy and Saxton (2012), captured by community frames in our work. In line with such previous work, we found that the use of action-infused social media messages (i.e., our motivational frame) was generally very low. At the same time, substantial heterogeneity emerged across unions, again in line with previous research on diverse communication repertoires. Importantly, we observed systematic differences by organizational structure: industrial unions—organizing across skill boundaries—tended to use diagnostic and prognostic frames more than the baseline, whereas craft unions—organized around specialized skill sets—used them less. This pattern resonates with their distinct organizational orientations, with industrial unions more often engaged in large-scale mobilization and bargaining.
Second, in the lead-up to elections, unions in won cases systematically used nearly all frames more frequently than in lost cases, especially diagnostic, prognostic, and community (RQ2). This finding matches the literature on innovative strategies and union revitalization (Turner 2005), suggesting that successful unions may engage in more versatile and intensive framing strategies. It also aligns with research linking injustice frames and solidarity appeals to mobilization outcomes (Benford and Snow 2000; Pasquier and Wood 2018).
Third, we found dynamic adjustments in the aftermath of elections: won cases more often showed decreases in prognostic and motivational framing compared to lost cases (RQ3). This pattern suggests a shift in priorities once the immediate goal of winning representation is secured, with communication recalibrated from the mobilization to the consolidation phase. Such recalibration aligns with broader theories on organizational life cycles of social movements, considering transitions from mobilization to institutionalization (Christiansen 2009; Tarrow 2022).
Taken together, these findings connect directly to longstanding debates about revitalization of labor unions. Scholars have argued that unions must adapt not only organizationally but also communicatively (Milkman 2013), experimenting with innovative repertoires (Juravich and Bronfenbrenner 1998; Voss and Sherman 2000). By systematically distinguishing between the communication practices around successful and unsuccessful election outcomes, our study suggests how strategic framing might be an important part of these adaptation processes. More broadly, we show the value of micro-level, comparative framing analysis for advancing our understanding of how movements strategize and evolve in the digital era.
Limitations, Implications and Future Work
While our study makes both methodological and substantive contributions, several limitations should be noted. First, our data is correlational, so observed associations between features and success cannot be interpreted causally. Second, the analysis is limited to Facebook—an important but partial element of unions’ communication strategies, which also involve other digital platforms and offline channels. Third, we focused on national and international level union accounts to ensure sufficient activity, which may obscure dynamics in smaller or more localized organizations. Fourth, we analyzed five discourse frames identified in prior work on social movements and civil society organizations. This enhances comparability but may overlook additional frames or nuances. Fifth, our pre- and post-election windows captured short-term dynamics but may miss longer-term adjustments. Finally, unions differ in size, resources, and political context, factors likely shaping communication in ways not fully accounted for here. We also acknowledge that systematic analyses of union communication could be misused by employers or political opponents, underscoring the need for ethical reflection in applying these methods.
Despite these limitations, the study has implications for multiple audiences. For labor unions, the findings can inform digital communication strategies, highlighting how diagnostic, prognostic, community, motivational, and engagement frames can be used in election campaigns. For political communication researchers, the study offers large-scale empirical evidence on how civil society organizations employ framing online. For organizations more broadly, our results suggest ways to refine messaging to balance mobilization, solidarity, and engagement objectives. Methodologically, the released classifier and dataset provide a scalable resource for studying digital discourse framing across movements.
Future work could extend this study in several directions. Expanding the frame taxonomy would allow more fine-grained analyses of communication strategies, including issue-specific frames. Examining union accounts at smaller, local levels—beyond the national and international accounts analyzed here—could reveal additional variation in messaging. Multi-platform analyses would provide a more holistic picture of unions’ communication ecosystems beyond Facebook. Incorporating a perspective on the targets of unions’ communication would also provide a broader understanding of strategic efforts, and complement existing literature suggesting ad hoc targeting as one valuable revitalization strategy (Earl 2019; Uba and Jansson 2021). Finally, incorporating audience responses through engagement metrics and comments would help link message production with resonance, offering deeper insight into the interplay between framing practices and collective action dynamics.
Related Work
NLRB Election Data
The National Labor Relations Board (NLRB) election data has been widely used to study the dynamics of unionization in the United States. A central line of research has focused on identifying predictors of union success. Early work by Maranto and Fiorito (1987) highlights union size, internal democracy, strike activity, and centralized decision-making as key determinants of certification outcomes. Extending this perspective, Ferguson (2008) builds a sequential model of organizing, tracing the process from election petitions to first contracts, and identifies union size and unfair labor practice charges as critical factors across stages.
A second strand of studies combines NLRB data with external sources to investigate how broader institutional and demographic contexts shape unionization. For example, Sojourner and Yang (2022) links NLRB elections with Occupational Safety and Health Administration data to show how certification influences the enforcement of workplace safety laws. Similarly, Ferguson (2015) integrates Equal Employment Opportunity Commission surveys to examine how unionization affects workplace composition, with attention to racial and social diversity. Bronfenbrenner (2005) takes a related approach by comparing bargaining units across gender compositions, assessing the role of company characteristics, union demographics, and organizing tactics.
Other contributions place NLRB election outcomes within broader structural or political-economic dynamics. Stepan-Norris and Southworth (2010) analyzes competition among AFL-CIO affiliates, showing how inter-union rivalry shapes union density when accounting for political and economic conditions. Schaller (2023) instead focuses on long-term union decline, attributing it to sectoral and regional shifts in economic activity. Finally, Ferguson, Dudley, and Soule (2018) connects labor organizing to wider social movements, introducing the concept of osmotic mobilization to demonstrate how protest activity in society at large correlates with support for unions in representation elections.
This literature demonstrates the versatility of NLRB election data for analyzing both the internal dynamics of organizing and its intersection with external institutional, demographic, and political factors. While existing studies have focused on predictors of election success, workplace outcomes, and long-run trends, our approach incorporates the digital domain and employs this data as an anchor to investigate transformations in digital communication.
Digital Communication of Labor and Social Movements
Labor Movement
With the large-scale adoption of digital technologies, labor unions have turned to online platforms as tools to engage with members, recruit new participants, and communicate with stakeholders. Scholars emphasize that unions’ strategies are shaped not only by external political contexts, but also by internal organizational choices (Martínez Lucio 2003). In this regard, Frege and Kelly (2004) identifies six strategic choices related to union revitalization, among which organizing—recruiting, representing, and retaining members, as well as expanding mobilization capacity—is particularly relevant in the digital context. Within these revitalization strategies, communication is central: discourse frames provide unions with a way to articulate grievances, propose solutions, and mobilize collective identities, thereby directly supporting organizing efforts. Building on this perspective, a consistent body of literature has linked framing and the labor movement.  Kelly (1998) adapts the collective action frames introduced by Snow, Benford et al. (1988) to the labor context and argues that successful organizing depends on workers interpreting grievances as unjust, identifying solutions, and believing collective action can be effective. Similarly, Voss and Sherman (2000) highlights the importance of identifying the right strategic repertoires of communication and engagement for the labor revitalization efforts to be successful. Other foundational work that followed highlights the importance of identifying the right frames for women and minorities (Bronfenbrenner 2005) and millenials (Milkman 2013).
Much of the literature on revitalization through digital media communication has focused on mobilization. Studies of the Fight for Fifteen movement, for example, use qualitative and mixed-methods analysis to demonstrate how digital media facilitate mobilizing efforts and illuminate the interplay between elements of collective action and the connective logic of grassroots initiatives (Pasquier and Wood 2018; Pasquier, Daudigeos, and Barros 2020; Frangi, Zhang, and Hebdon 2020). Similarly, Houghton and Hodder (2021) focuses on mobilizing frames across platforms, coding social media posts in terms of framing, attribution, and calls to action. Beyond mobilization, research has also examined how digital networks support organizational purposes, such as building collective voices prior to mobilization (Lazar, Ribak, and Davidson 2020), or reaching marginalized groups otherwise excluded from traditional organizing channels (Wood 2015; Thornthwaite, Balnave, and Barnes 2018). The effectiveness of communication strategies is also contingent on other important factors, such as organizational and leadership structures, as demonstrated in the comparative analysis of 2018 teachers’ strikes in Oklahoma and Arizona (Blanc 2022).
Despite these insights, evidence suggests that unions are not fully exploiting the potential of digital media.  Hodder and Houghton (2020) finds that online campaigning and news reporting remain dominant uses, leaving recruitment opportunities largely untapped. Likewise, Carneiro and Costa (2022) documents outdated strategies centered on one-way information sharing, rather than fostering interactive or targeted engagement. More promising results emerge when unions tailor their communication to specific worker groups (Jansson and Uba 2019) and account for the diversity of audiences with distinct communication needs (Panagiotopoulos 2021). These targeted strategies appear to be a necessary step towards building more effective digital communication.
Beyond the Labor Movement
Beyond unions, a large body of research examines how social movements and civil society organizations use digital platforms to structure collective action and mobilize supporters.
One line of work emphasizes the role of mobilization frames in online environments. Building on framing theory, scholars show how social media enable the rapid diffusion of diagnostic, prognostic, and motivational frames, which are crucial for lowering barriers to participation (Benford and Snow 2000; Fernández-Zubieta et al. 2023; Mendelsohn et al. 2024; Pera and Aiello 2025). A second perspective highlights the shift from the logic of collective action toward the logic of connective action (Bennett and Segerberg 2012). While traditional mobilization relies on formal organizations and collective identities, connective action underscores personalized, digitally mediated participation that can occur without strong organizational structures. This framework has been widely applied to cases such as the Arab Spring (Steinert-Threlkeld 2017) and climate activism (Segerberg and Bennett 2011). Finally, research on digital affordances investigates how platform features—such as sharing content, meta-voicing, and multimedia storytelling—shape the opportunities and constraints of collective action (Zheng and Yu 2016; Sæbø, Federici, and Braccini 2020).
While prior research has provided valuable insights into social media mobilization and organizational communication, many studies rely on qualitative analyses or selected use cases. In contrast, our study leverages a large-scale computational approach to systematically examine union social media communication during times of representation elections at the micro-level of individual messages. Moreover, by linking communication patterns to NLRB election outcomes, we explore how online framing relates to tangible organizational success, providing empirical evidence on the functional value of digital strategies that complements existing theoretical and qualitative work.
Acknowledgments
We acknowledge the support from the Carlsberg Foundation through the COCOONS project (CF21-0432). The funder had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.
Appendix A Frames Annotation
We annotated Facebook posts for the presence of five discourse frames: diagnostic, prognostic, motivational, community, and engagement. In addition, given the close ties between labor unions and political events (e.g., presidential elections, support for congressional candidates), we also annotated posts with an auxiliary label: political endorsement. This label captures instances in which unions explicitly support political actors or candidates.
The complete annotation guidelines are provided in the supplementary material (codebook.pdf).
Appendix B Modeling and Computational Details
We fine-tuned a RoBERTa-base model to capture the characteristics of unions’ Facebook communications. For this fine-tuning of the language model, we used 5 epochs with a batch size of 128, keeping all other parameters at their default values. Training was performed on a Tesla V100 GPU and required about 30 minutes.
For the multi-label classification task, we trained the RoBERTa model for 30 epochs with a learning rate of $4\times 10^{-5}$ and a batch size of 32. To address class imbalance, we applied class weights computed as
$\frac{\text{total samples}}{2\times[n_{\text{class 0}},\ldots,n_{\text{class n}}]}.$
All other parameters were left at default settings. Training required a few minutes on a Tesla V100.
The multi-label classification setup included the five main discourse frames and the additional political endorsement label. However, given the very limited presence of political endorsement in the annotated dataset (3% of posts across training and test sets) and in the full Facebook dataset (2.5% of posts), we did not retain this label in the main analysis. We therefore focus exclusively on the five core discourse frames in the paper.
B.1 Licenses
The RoBERTa base model used for building the multi-label communication frames classifier is licensed under a MIT license.
Appendix C Robustness Checks
C.1 Window Size for Rolling Aggregation
The usage of frames is robust to the choice of the window size for rolling aggregation: despite minor changes, differences between won and lost representation cases are stable both pre- (cf. Figure C1) and post-elections (cf. Figure C2). For all plots, we balanced the number of wins and losses per union.
((a)) $n_{days}=5$
((b)) $n_{days}=3$
Figure C1: Usage of frames leading to elections, for lost and won cases. Results are reported in terms of mean and 95% confidence interval. Values on the y-axis refer to detrended scores. The $\star$ sign refers to the significance of the test for difference in means between the losing and winning: $\star$ indicates significance at $\alpha=0.05$, $\star\star$ at $\alpha=0.01$, and $\star\star\star$ at $\alpha=0.001$.
((a)) $n_{days}=5$
((b)) $n_{days}=3$
Figure C2: Usage of frames post-elections, for lost and won cases. Results are reported in terms of mean and 95% confidence interval. Values on the y-axis refer to detrended scores. The $\star$ sign refers to the significance of the test for difference in means between the losing and winning: $\star$ indicates significance at $\alpha=0.05$, $\star\star$ at $\alpha=0.01$, and $\star\star\star$ at $\alpha=0.001$.
C.2 Imbalance for Number of Cases per Union
Frames Usage Leading to Elections
To assess robustness to the choice of random seed for sampling when balancing the number of wins and losses per union, we repeated the sampling across 20 random seeds. For each seed and frame, we performed independent t-tests comparing the difference in means between lost and won cases. Table C1 summarizes the results, reporting the percentage of seeds (over the 20 in total) for which the difference is significant at standard thresholds, along with median $t$-statistics, $p$-values, and ranges. Differences in all frames remain consistently significant.
Statistic Diagnostic Prognostic Motivational Community Engagement
% $p<0.05$ 100% 100% 100% 100% 100%
% $p<0.01$ 100% 100% 100% 100% 100%
% $p<0.001$ 100% 100% 100% 100% 100%
Median $t$ -11.11 -15.28 -7.17 -5.49 5.42
Median $p$ $2.2\times 10^{-28}$ $1.7\times 10^{-51}$ $8.6\times 10^{-13}$ $4.2\times 10^{-8}$ $6.1\times 10^{-8}$
Range $t$ [-12.53, -10.29] [-16.52, -14.43] [-8.03, -6.85] [-6.38, -4.11] [4.51, 6.44]
Table C1: t-test results for the difference in means between lost and won cases across 20 random seeds for balancing wins and losses per union. Reported are the percentage of tests below standard significance thresholds, the median $t$-statistic and $p$-value, and the range of $t$-statistics.
Next, we aimed at assessing the robustness to differences in the number of representation cases per union. We limited the total cases per union to the 90th percentile of the distribution (considering winning and losing outcomes jointly), and repeated the analyses across 20 random seeds. We first balanced the number of wins and losses per union for each seed. Then, for each seed and frame, we performed independent t-tests comparing the difference in means between lost and won cases. Table C2 summarizes the results, reporting the percentage of seeds (over the 20 in total) for which the difference is significant at standard thresholds, along with median $t$-statistics, $p$-values, and ranges. Differences in the diagnostic, prognostic, and community frames remain consistently significant, while motivational and engagement frames are not robust to the down-sampling procedure.
Statistic Diagnostic Prognostic Motivational Community Engagement
% $p<0.05$ 100% 100% 85% 100% 95%
% $p<0.01$ 100% 100% 25% 100% 65%
% $p<0.001$ 100% 100% 0% 90% 10%
Median $t$ -6.16 -7.95 -2.34 -3.66 2.66
Median $p$ $8.1\times 10^{-10}$ $2.5\times 10^{-15}$ $1.9\times 10^{-2}$ $2.5\times 10^{-4}$ $7.9\times 10^{-3}$
Range $t$ [-6.79, -5.44] [-8.46, -7.31] [-2.93, -1.42] [-4.31, -3.17] [1.85, 3.71]
Table C2: t-test results for the difference in means for lost and won cases across 20 random seeds. Reported are the percentage of tests below standard significance thresholds, the median $t$-statistic and $p$-value, and the range of $t$-statistics.
Frames Usage Pre- and Post-Elections
Given that we balanced the number of wins and losses per union, we verified the robustness of results to the choice of random seed for sampling. As such, we repeated the balancing procedure with 20 random seeds. For each combination of pattern, frame and seed, we then computed the offset between lost and won cases (losing minus winning) and measured the significance of such an offset. In the end, we considered the average offset value across seeds for each pattern and frame combination and highlighted the offset as significant if 80% of the seeds showed a significant offset for that combination. Figure C3 summarizes these robustness results. The most relevant findings are robust: communication in won cases showed a higher frequency of decrease in prognostic and motivational frames after the election, and communication in lost cases was characterized by greater stability in the use of the motivational frame.
Figure C3: Usage of frames before and after the event, comparing lost and won cases. We report raw counts of cases assigned to each cell for losses (L) and wins (W). The $\star$ sign indicates a significant difference between outcome types at $\alpha=0.05$.
To account for potential imbalance in the number of representation cases per union, we limited the total cases per union to the 90th percentile, repeating the analyses across 20 random seeds. For each seed and frame, post-/pre-election changes were classified into three patterns: decrease, stable, or increase. For each seed, we first balanced the number of wins and losses per union. Then, for each combination of pattern, frame and seed, we computed the offset between lost and won cases (losing minus winning) and measured the significance of such an offset. In the end, we considered the average offset value across seeds for each pattern and frame combination and highlighted the offset as significant if 80% of the seeds showed a significant offset for that combination. Figure C4 summarizes these robustness results. While not all patterns remain significant under this procedure, the most relevant findings are robust: communication in won cases showed a higher frequency of decrease in prognostic and motivational frames after the election.
Figure C4: Usage of frames before and after the event, comparing lost and won cases. We report raw counts of cases assigned to each cell for losses (L) and wins (W). The $\star$ sign indicates a significant difference between outcome types at $\alpha=0.05$.
Appendix D Details on Labor Unions
Table D1 reports the acronyms and full names for national-level unions considered in the study. Figure D1 shows the distribution of the number of Facebook posts per labor union in the 2015-2024 period, highlighting some skewness.
Acronym Expanded Name Union Structure
RWDSU Retail, Wholesale and Department Store Union Industrial
UAW United Automobile Workers Industrial
IBEW International Brotherhood of Electrical Workers Craft
AFT American Federation of Teachers Industrial
NATCA National Air Traffic Controllers Association Craft
IFPTE International Federation of Professional and Technical Engineers Industrial
AFSCME American Federation of State, County and Municipal Employees Industrial
NNU National Nurses United Industrial
IUPAT International Union of Painters and Allied Trades Craft
Ironworkers International Association of Bridge, Structural, Ornamental and Reinforcing Iron Workers Craft
NEA National Education Association Industrial
UMWA United Mine Workers of America Industrial
AFGE American Federation of Government Employees Industrial
IATSE International Alliance of Theatrical Stage Employees Craft
UFCW United Food and Commercial Workers Industrial
SIU Seafarers International Union of North America Industrial
SEIU Service Employees International Union Industrial
SAG-AFTRA Screen Actors Guild – American Federation of Television and Radio Artists Craft
BCTGM Bakery, Confectionery, Tobacco Workers and Grain Millers Industrial
APWU American Postal Workers Union Industrial
UWUA Utility Workers Union of America Industrial
BAC Bricklayers and Allied Craftworkers Craft
SMART Sheet Metal, Air, Rail and Transportation Workers Craft
AFM American Federation of Musicians Craft
CWA Communications Workers of America Industrial
USW United Steelworkers Industrial
UNITE HERE Unite Here Union Industrial
IAM International Association of Machinists and Aerospace Workers Industrial
IBT International Brotherhood of Teamsters Industrial
UA United Association of Journeymen and Apprentices of the Plumbing and Pipe Fitting Industry Craft
TWU Transport Workers Union of America Industrial
IAFF International Association of Fire Fighters Craft
ATU Amalgamated Transit Union Industrial
IBB International Brotherhood of Boilermakers, Iron Ship Builders, Blacksmiths, Forgers and Helpers Craft
Roofers United Union of Roofers, Waterproofers and Allied Workers Craft
LIUNA Laborers’ International Union of North America Craft
WGAE Writers Guild of America East Craft
OPEIU Office and Professional Employees International Union Industrial
CSEA Civil Service Employees Association Industrial
MEBA Marine Engineers’ Beneficial Association Craft
Table D1: Union acronyms, expanded names, and organizational structure (Craft vs. Industrial).
Figure D1: Number of Facebook posts by labor union.